find the length and width of a square
area=144 square centimeters
perimeter=48 centimeters

Answer by checkley75(3666)

perimeter of a square=4*side
48=4*s
s=48/4
s=12 answer for the length & width of this square.
proof:
area=s^2
144=12^2
144=144